# Channel Measurement, Modeling, and Simulation for 6G: A Survey and Tutorial

### v3

@@ VIII-A Development of Channel Simulators for 6G
In each generation, a few channel simulators were developed and played important roles in performance evaluations of communication technologies. In the 4G era, WINNER II provided the WINNER II channel models and corresponding implementations for link and system-level simulations [66], including generic and clustered delay line models defined for different propagation scenarios. In the 5G era, several organizations have developed simulators for link and system-level channel simulations. In 2014, the Quasi Deterministic Radio Channel Generator (QuaDriGa) was released by Fraunhofer Heinrich-Hertz-Institute [255]. In 2017, New York University (NYU) released a channel simulator named NYUSIM [256] based on the statistical spatial channel model proposed by NYU. In 2018, IMT-2020 CM-BUPT [257] was launched as an implementation of the widely recognized 5G standard GBSM channel model defined in ITU-R M.2412 [42] and 3GPP TR38.901 [43]. In recent years, with the development of 6G research, a series of channel simulators were released to capture channel characteristics of 6G technologies. In 2020, SimRIS was released [258], providing a propagation channel for RIS-assisted communication research in various indoor and outdoor environments. In 2021, researchers from the Higher Institute for Applied Sciences and Technology (HIAST) published the TeraMIMO channel simulator for wideband MIMO THz communications [259]. Besides, in 2023, the NYUSIM was updated to the 4.0 version which supports simulations of THz frequency bands.
Figure 21: Simulation framework for the standard GBSM channel model.
Figure 22: Simulation framework of BUPTCMCCCMG-IMT2030. The blue parts in Module II are simulation steps related to the basic simulation, and the green parts are related to the modifications for 6G.
@@
To meet the requirements of channel simulations supporting higher frequency bands, larger-scale antenna array, ISAC, RIS, and SAGIN, our team developed a 6G channel simulator called BUPTCMCCCMG-IMT2030 based on the framework of the 5G standard GBSM channel model. It supports the generation of CIRs for various scenarios, including urban macrocell (UMa), UMi, rural macrocell (RMa), InH, and SAGIN. New channel features and parameters are also integrated into the platform to support 6G key technologies, such as THz, E-MIMO, ISAC, and RIS.
VIII-B Basic Simulation Framework of BUPTCMCCCMG-IMT2030
@@
BUPTCMCCCMG-IMT2030 is implemented based on the procedure-oriented framework of the 5G standard GBSM channel model with a series of extensions for 6G features. As shown in Fig. 21, the framework of the basic simulation procedure is briefly summarized as follows[43].
TABLE VI: Detail simulation procedure for the astandard GBSM channel model
Steps Descriptions
Step 1 Set the scenario, network layout, and antenna parameters.
Step 2 Assign the propagation condition (LOS/NLOS) according to the network layout and LOS possibility.
Step 3 Calculate the path loss. The path loss models for the calculation are different for different scenarios, frequencies, and propagation conditions.
Step 4 Calculate the LSP, such as DS, AS, K-factor, and shadow fading. The parameters needed for the calculation are correlated to the scenario, frequency, and propagation condition.
Step 5 Generate cluster delays using DS in Step 4.
Step 6 Generate cluster powers based on delays in Step 5.
Step 7–8 Generate arrival and departure angles for clusters and rays based on the AS obtained in Step 4.
Step 9 Generate the XPRs.
Step 10 Draw random initial phases of each ray.
Step 11 Generate the channel coefficient.
Step 12 Apply path loss and shadowing.
@@ VIII-C Software Structure and Functions
A software platform structure of BUPTCMCCCMG-IMT2030 is illustrated in Fig. 22. The software consists of three modules: configuration, simulation, and analysis. In Module I, the configuration module, system parameters, such as scenario, center frequency, and bandwidth, is defined. In Module II, the simulation module, BUPTCMCCCMG-IMT2030, performs a simulation and generates a series of channel parameters. In Module III, the analysis module, simulation results are analyzed and shown in numerical and graphical formats. A stepwise description of the simulation is given as follows.
VIII-C1 Module I–Configuration
In the configuration module, there are two types of configurations: general configurations and 6G feature configurations. General configurations include basic system parameters for a complete simulation, including the scenario, center frequency, bandwidth, positions of BS and UE, antenna type and layout, and mobility of UE. 6G feature configurations are a series of parameters for new channel features for 6G. Various 6G channel features are supported, including an extension of fast-fading parameters and sparsity characteristics for the THz channel, near-field and SnS characteristics for the E-MIMO channel, and other features for channels, i.e., SAGIN, ISAC, and RIS.
@@ VIII-C1 Module I–Configuration
THz
Step 3,
Step 4,
Step 11
Step 3: Extend the path loss model for the THz frequency band; Step 4: Extend parameters for the THz frequency band; Step 11: Apply the sparsity to channel coefficients.
E-MIMO
(Near-field
effect, SnS)
Step 1,
Step 11
@@
Step 1: Set the target locations and geometric parameters; Step 3: Add the path loss calculation of the sensing channels; Step 4: Add the correlation between the sensing and communication channels; Step 5: Add the calculation of delays for the sensing channels; Step 7-8: Add the calculation of ray angles, RCS, and clutter for the sensing channels; Step 9-12: Add the calculation of the sensing CIR.
@@
(Cascade channel)
Step 1-11 Step 1: Set the configuration of RIS; Step 2-10: Propagation condition, path loss, LSP, and SSP are generated for both                  Tx-RIS and RIS-Rx links; Step 11: The cascade channel coefficient of Tx-RIS-Rx is calculated considering the               radiation pattern of RIS.
@@
Step 1-4 Step 1: Set the configuration of BS and UE in the ECEF coordinate system; Step 2-4: Models and parameters for the SAGIN scenarios are used.
VIII-C2 Module II–Simulation
@@
The 6G extensions segment includes five modifications for the 6G new channel features to the basic simulation segment. A summary of the modified steps and descriptions is shown in Table VII.
@@
    For the THz channel simulation, steps 3, 4, and 11 are modified. In step 3, the path loss model is extended to the THz frequency band. In step 4, the parameters for generating multipaths are extended to the THz frequency band. In step 11, the model described in [149] is applied to introduce sparsity into the THz channel.
  • Modifications for the E-MIMO channel simulation.
    For the E-MIMO channel simulation, steps 1 and 11 are modified. In step 1, the size of the station region, which is the number of antenna elements within the station region, is configured. In step 11, the spherical wave and the SnS characteristic are applied to calculate the channel coefficient. The phase evolution of the spherical wave is calculated based on the distance between antenna elements and scatterers, which is estimated according to the angles and delays of clusters. The SnS characteristic is simulated by applying the visibility variation of clusters based on a two-state Markov chain.
  • Modifications for the ISAC channel simulation.
    For the ISAC channel simulation, a series of steps are modified. The original steps are used to generate the communication channels, and the modifications are used to generate sensing channels. In step 1, besides the layout of the BS and UE, the location and geometric parameters of the sensing targets are configured. In step 3, the path loss of the sensing channels is calculated according to the configurations of the sensing targets. In step 4, the correlation between the sensing and communication channels is applied. In step 5, the delays of the sensing channels are calculated. In steps 7-8, a series of SSPs of the sensing channels, such as ray angles, RCS, and clutter, are calculated. In steps 9-12, the channel coefficients of both sensing channels and communication channels are calculated.
@@
    For RIS-assisted channel simulation, BUPTCMCCCMG-IMT2030 has modified a series of steps based on the 5G standard channel model. In the RIS-assisted channel, the channel is divided into two links, Tx-RIS and RIS-Rx, and parameters are generated for them and calculated together. In step 1, the RIS is configured apart from the layout of the BS and UE. In steps 2-10, the propagation condition (LOS/NLOS), path loss, LSPs, and SSPs are generated for both Tx-RIS and RIS-Rx links. In step 11, the cascade channel coefficients of the RIS-assisted channel are calculated by combining the channel coefficients of the Tx-RIS and RIS-Rx links and applying the radiation pattern of the RIS.
@@
    For the SAGIN channel simulation, the channel model defined in 3GPP TR 38.811 [254] is implemented and integrated into BUPTCMCCCMG-IMT2030, including modifications to the coordinate system, models, and parameters. In step 1, the ECEF coordinate system is used to locate the BS and UE. In step 2, the propagation conditions (LOS/NLOS) are assigned using a different method. In step 3, the path loss model is modified as described in VII-C. In step 4, the fast-fading parameters for SAGIN are used, which are related to both the frequency and elevation angle.
@@
THz
Scenario: UMi-street canyon; LOS status: LOS
Center frequency: 132 GHz; Bandwidth: 1.2 GHz
Position (x,y,z): BS:[0,0,3] (m); UE:[10,10,3] (m)
Antenna: BS: Single antenna; UE: Single antenna
Intracluster K-factor: 17.98 dB
E-MIMO
Scenario: UMi-street canyon; LOS status: LOS
Center frequency: 28 GHz; Bandwidth: 200 MHz
Position (x,y,z): BS:[0,0,3] (m); UE:[10,10,3] (m)
Antenna: BS: 256 elements ULA; UE: Single antenna
Stationary region size: 16 elements
@@
RIS size: 32$\times$32 elements
@@
ASA: [$10^{\circ}$, $5^{\circ}$, and $1^{\circ}$]
@@
VIII-C3 Module III–Analysis
In Module III, a simple analysis of numerical and graphical results based on the generated channel coefficients is provided, including delay spread, angular spread, the reference signal receive power (RSRP), Gini index (used for sparsity analysis), network layout, and CIR of a single link.
@@ VIII-D Applications
In this section, some application examples of the typical 6G technologies are provided. The main simulation configurations for different 6G cases are listed in Table VIII.
VIII-D1 6G Feature 1–THz Frequency Band
@@
Using the configurations in Table VIII, the CIR results of the generated channel at 132 GHz in the THz band were compared with those at 28 GHz in the mmWave band, as shown in Fig. 23. It can be found that the THz channel has 8 clusters and the mmWave channel has 16 clusters. Fewer multipaths are present in the THz band than in the mmWave band, which indicates that the THz channel has a more sparse structure than the mmWave channel. In addition, the delay spread of the THz channel (4.27 ns),is much lower than that of the mmWave channel (30.94 ns) according to Fig. 23. The multipaths of the THz channel are concentrated in a narrower delay range, which means that multipaths in the THz channel are concentrated in a small time period.
VIII-D2 6G Feature 2–E-MIMO
Based on the configurations in Table VIII, the absolute cross-correlation coefficient of CIRs between the reference element and other elements in comparison with a channel generated using the standard channel model is shown in Fig. 24. With the increase in spacing between two elements, the cross-correlation coefficient of the standard channel model remains stable, but that of the E-MIMO channel continues to decrease, indicating that the correlation of the two elements becomes weaker. In addition, the appearance or disappearance of clusters and the evolution of the phase also occur. The appearance or disappearance of clusters results in a sudden change in the cross-correlation coefficients, including an increase in the appearance and a decrease in the disappearance. When no cluster appears or disappears, the evolution of the phase results in a change in the cross-correlation coefficients.
VIII-D3 6G Feature 3–ISAC
Figure 25: Simulation results of the SD ($\text{SD}_{\text{s}}$) in the sensing channels under different numbers of shared clusters ($N_{0}$) when total clusters ($N$) are 15.
Using the configurations in Table VIII, we can generate the shared and nonshared clusters of the proposed ISAC channel model in (17). The CDF of the SD metric in the sensing channels, which is the PR of the shared clusters ($N_{0}$) to the total clusters ($N$) in the ISAC channel, under different numbers of shared clusters is shown in Fig. 25. A higher SD indicates a greater potential to realize the mutual auxiliary functions of communication and sensing. As $N_{0}$ increases, $\text{SD}_{\text{s}}$ also increases. By controlling the appropriate $N_{0}$ of the proposed channel model, the desired SD value can be achieved, facilitating flexible simulation implementation of ISAC channels.
@@
Using the configuration in Table VIII, the impact of channel angular spread on the performance of RIS-assisted communication is illustrated in Fig. 26. With the increase in azimuth angle spread of arrival (ASA) in the Tx-RIS channel, the performance of the RIS cascade communication link will worsen. This is because the greater the angular spread, the lower the energy in the specified incident direction. Additionally, the impact of the angle-dependent characteristics of RIS reflection coefficients on RIS-assisted communication, as mentioned in Section VI-B1, has been considered in the simulation. In the legends, ”Non-ideal modulation” represents the simulation results when considering this angle-dependent characteristic. It can be observed that the non-ideal modulation results in approximately 0.5, 0.7, and 1 dB of attenuation under the three angle spreads.
@@
Using the configurations in Table VIII, the path loss of SG channels with different channel configurations is illustrated in Fig. 27. The path loss of SG channels is affected by the height of satellites, the elevation angle of the link, and the center frequencies. The increase in satellite height results in an increase in propagation distance, which makes the path loss larger. Furthermore, the path loss increases when a higher frequency band is used. Additionally, the elevation angle of the link can affect the path loss in SG scenarios; when the elevation angles increase, the path loss becomes larger.

### v4

@@ VIII-A Development of Channel Simulators for 6G
In each wireless generation, a few channel simulators were developed and played important roles in performance evaluations of communication technologies. In the 4G era, WINNER II provided channel models and corresponding implementations for link- and system-level simulations [66], including generic and clustered delay line models defined for different propagation scenarios. In the 5G era, several organizations have developed simulators for link- and system-level channel simulations. In 2014, the Quasi Deterministic Radio Channel Generator (QuaDriGa) was released by the Fraunhofer Heinrich-Hertz Institute [300]. In 2017, New York University (NYU) released a channel simulator named NYUSIM [301], which is based on the statistical spatial channel model proposed by NYU. In 2018, IMT-2020 CM-BUPT [302] was launched as an implementation of the widely recognized 5G standard GBSM channel model defined in ITU-R M.2412 [46] and 3GPP TR 38.901 [47]. In recent years, with the development of 6G research, a series of channel simulators have been released to capture the channel characteristics of 6G technologies. In 2020, SimRIS was released [303], providing a propagation channel for RIS-assisted communication research in various indoor and outdoor environments. In 2021, researchers from the Higher Institute for Applied Sciences and Technology (HIAST) published the TeraMIMO channel simulator for wideband MIMO THz communications [304]. In addition, in 2023, the NYUSIM was updated to version 4.0, which supports simulations of THz frequency bands. In response to potential 6G technologies, including ISAC, XL-MIMO, THz, and RIS, research institutions and universities worldwide have developed various simulation platforms. The team from New York University compiled a comprehensive global survey of simulators in 2023 [305], encompassing channel simulators, link-level simulators, system-level simulators, and network simulators. In this work, we specifically focus on conducting a systematic review of 6G channel simulators worldwide, as summarized in Table VI. The analysis reveals that the majority of channel simulators target THz band characterization, whereas a minority address specific domains. For example, NirvaWave specializes in near-field channel modeling, and SimRIS focuses on RIS channel emulation. Notably, BUPTCMCCCMG-IMT2030 stands out as one of the few simulators capable of supporting multiple emerging technologies simultaneously.
Figure 26: Simulation flowchart for the BUPTCMCCCMG-IMT2030 based on the E-GBSM model.
Figure 27: Principle of the BUPTCMCCCMG-IMT2030 simulator based on the E-GBSM model [98].
Figure 28: Simulation framework of BUPTCMCCCMG-IMT2030.
@@
To meet the requirements of channel simulations supporting higher frequency bands, larger-scale antenna array, ISAC, RIS, and SAGIN, our team developed a 6G channel simulator called BUPTCMCCCMG-IMT2030 on the basis of the framework of the 5G standard GBSM channel model. It supports the generation of CIRs for various scenarios, including urban macrocells (UMa), UMi, rural macrocells (RMa), InH, and SAGIN. New channel features and parameters are also integrated into the platform to support 6G key technologies, such as ISAC, XL-MIMO, THz, and RIS.
VIII-B Simulation principle of BUPTCMCCCMG-IMT2030
@@
BUPTCMCCCMG-IMT2030 is implemented on the basis of the procedure-oriented framework of the 5G standard GBSM channel model, with a series of extensions for 6G features. As shown in Fig. 26, the framework of the basic simulation procedure is briefly summarized as follows [47]. To enable these extensions in a systematic way, the platform relies on the extended GBSM (E-GBSM) principle as its theoretical foundation. This principle provides a unified mechanism to incorporate new channel characteristics beyond the 5G model.
The E-GBSM principle has been detailed in [98]. In this paper, we take several representative 6G features—such as near-field effects, SnS, channel concatenation, and target RCS—as examples to illustrate how these characteristics can be incorporated into a unified expression under the E-GBSM framework. The basic principle is depicted in Fig. 27.
The first equation, $\mathbf{r}_{p,x}$, represents the non-concatenated channel response in the angle and delay domains, such as for Tx–Target, Target–Rx, or standard Tx–Rx communication links. The components highlighted in red boxes indicate the extensions introduced for various 6G channel features. For example, $S_{p,n_{t}}^{\mathrm{Tx}}$ describes the SnS behavior at the Tx antenna array (i.e., visibility of clusters), as defined in Eq. 11. $A_{p,n_{t},m_{t}}^{\mathrm{Tx}}$ models the near-field effect through a second-order phase correction, with its expression provided in Eq. 10. When modeling high-frequency sparsity, the parameter $I_{n_{t},m_{t}}$ (intra-cluster K-factor, ICK) can be introduced to redistribute power among paths within a cluster, concentrating energy around the LOS component [51]. Additionally, if the sharing feature in the ISAC channel is considered, $N_{s}$ denotes the number of shared scatterers simultaneously contributing to both communication and sensing links, as described in earlier sections.
The second equation describes the time-domain representation of a concatenated channel, using the Tx–Target–Rx link as an example. Here, the amplitude gain whether representing a RIS radiation pattern or target RCS is denoted by $F^{\mathrm{Tar}}$, and the polarization interaction is modeled via the matrix $\mathrm{CPM}^{\mathrm{Tar}}$. The $_{\tau}$ symbol denotes a delay-domain convolution operation at the small-scale level, indicating that the overall Tx–Target–Rx channel is constructed by convolving the two subchannels, Tx–Target and Target–Rx, in the delay domain.
@@ VIII-C Software Structure and Functions
The software platform structure of BUPTCMCCCMG-IMT2030 is illustrated in Fig. 28. The software consists of three modules: configuration, simulation, and analysis. In Module I, the configuration module and system parameters, such as the scenario, center frequency, and bandwidth, are defined. In Module II, the simulation module, BUPTCMCCCMG-IMT2030, performs a simulation and generates a series of channel parameters. In Module III, the analysis module, simulation results are analyzed and presented in numerical and graphical formats. A stepwise description of the simulation is given below.
VIII-C1 Module I: Configuration
In the configuration module, there are two types of configurations: general configurations and 6G feature configurations. The general configurations include basic system parameters for a complete simulation, including the scenario, center frequency, bandwidth, positions of the BS and UE, antenna type and layout, and mobility of the UE. 6G feature configurations are a series of parameters for new channel features for 6G. Various 6G channel features are supported, including an extension of fast-fading parameters and sparsity characteristics for the THz channel; near-field and SnS characteristics for the XL-MIMO channel; and other features for channels, i.e., SAGIN, ISAC, and RIS.
TABLE VI: A comparative analysis of the channel simulators for 6G
Channel Simulator Developer Channel Simulation Capability Release Time
BUPTCMCCCMG-IMT2030 ARTT Lab of BUPT A channel model based on measurements ranging from 0.5 to 330 GHz [17], supporting 6G new features and technologies such as ISAC, near field, SnS, RIS, NTN, etc. The simulator is fully compatible with standard channel models such as ITU-R M.2412, 3GPP TR 38.900/901, TR 36.777, TR 36.873 and TR 37.885, and has been calibrated based on TR 38.901 as the baseline. The developer was fully involved in the 3GPP R19 ISAC and 7–24 GHz channel modeling agenda. The simulator possesses the latest R19 channel simulation capabilities. June 2023
QuaDRiGa v2.8.1 Fraunhofer Heinrich Hertz Institute (HHI) [300] Compatible with 3GPP TR 36.873, TR 37.885, TR 38.901 TDL, CDL channel models; compatible with mmMAGIC channel model, satellite channel modeling, and supports spatial consistency channel modeling. December 2023
NYUSIM v4.0 NYU WIRELESS [305] Based on measured statistical channel models, supporting UMi, UMa, RMa, InH, InF scenarios simulation from 0.5 to 150 GHz. June 2023
Thor Simulator Horizon 2020 Joint EU-Japan project [306] A simulator for 300 GHz backhaul links. June 2022
KUCG v1.0.1 Kyoto University [307] Used to generate channel impulse responses for the 60 GHz (millimeter wave), 95 GHz and 105 GHz (sub-THz) frequency bands. These responses are based on actual channel measurements conducted by the Orihara Laboratory of Kyoto University. May 2024
SimRIS v2.0.1 Koç University [300] Can be used for channel modeling in RIS-assisted MIMO systems, with adjustable working frequency, terminal position, number of RIS components, and environment. August 2023
NirvaWave Princeton University [308] A novel near-field channel simulator, built on scalar diffraction theory and Fourier principles, that precisely captures the propagation evolution of arbitrary user defined transmit electromagnetic (EM) signals in complex user defined wireless mediums. September 2024
DeepMIMO v2 Wireless Intelligence Lab REMCOM [309] Based on Remcom 3D ray tracing and parameter set generation to create large-scale MIMO datasets. 2021
BUPTCMCC-6G-DataAI+ ARTT Lab of BUPT [310] The BUPTCMCC-6G-DataAI+ employs a generative data construction method, generating up to 300 million sets of channel data. It further expands into new scenarios, such as maritime scenario, ultra-high-speed mobility, multiple-perspective sensing, industrial internet, and indoor RIS, and introduces a new intermediate frequency band of 7–24 GHz. This supports the requirements of air interface AI algorithms, including CSI compressed feedback, occlusion and beam prediction, sensing positioning, air interface resource management, and network planning optimization. September 2024
@@ VIII-C1 Module I: Configuration
Step 1: Set the target locations and geometric parameters; Step 3: Add the path loss calculation of the sensing channels; Step 4: Add the correlation between the sensing and communication channels; Step 5: Add the calculation of delays for the sensing channels; Steps 7-8: Add the calculation of ray angles, RCS, and clutter for the sensing channels; Steps 9-12: Add the calculation of the sensing CIR.
XL-MIMO
(Near-field
effect, SnS)
Step 1,
Step 11
Multi-Frequency
Step 3,
Step 4,
Step 11
Step 3: Extend the path loss model for the multi-frequency band; Step 4: Extend parameters for the multi-frequency band; Step 11: Apply selectively the sparsity to channel coefficients.
@@
(Concatenated channel)
Steps 1-11 Step 1: Set the configuration of RIS; Steps 2-10: Propagation condition, path loss, LSP, and SSP are generated for both         Tx-RIS and RIS-Rx links; Step 11: The concatenated channel coefficient of Tx-RIS-Rx is calculated considering       the radiation pattern of RIS.
@@
Step 1-4 Step 1: Set the configuration of BS and UE in the ECEF coordinate system; Steps 2-4: Models and parameters for the SAGIN scenarios are used.
VIII-C2 Module II: Simulation
@@
The 6G extension segment includes five modifications for the new 6G channel features to the basic simulation segment. A summary of the modified steps and descriptions is shown in Table VII.
  • Modifications for the ISAC channel simulation.
    For the ISAC channel simulation, a series of steps are modified. The original steps are used to generate the communication channels, and the modifications are used to generate sensing channels. In step 1, in addition to the layout of the BS and UE, the location and geometric parameters of the sensing targets are configured. In step 3, the path loss of the sensing channels is calculated according to the configurations of the sensing targets. In step 4, the correlation between the sensing and communication channels is applied. In step 5, the delays of the sensing channels are calculated. In steps 7-8, a series of SSPs of the sensing channels, such as ray angles, RCS, and clutter, are calculated. In steps 9-12, the channel coefficients of both the sensing channels and the communication channels are calculated. The detailed modifications and implementation procedures for ISAC channel simulation can be found in [311, 312].
  • Modifications for the XL-MIMO channel simulation.
    For the XL-MIMO channel simulation, steps 1 and 11 are modified. In step 1, the size of the station region, which is the number of antenna elements within the station region, is configured. In step 11, the spherical wave and SnS characteristics are applied to calculate the channel coefficient. The phase evolution of the spherical wave is calculated on the basis of the distance between the antenna elements and scatterers, which is estimated according to the angles and delays of the clusters. The SnS characteristic is simulated by applying the visibility variation of clusters on the basis of a two-state Markov chain. The simulation implementation of the XL-MIMO channel is shown in reference [313].
@@
    For the THz channel simulation, steps 3, 4, and 11 are modified. In step 3, the path loss model is extended to the THz frequency band. In step 4, the parameters for generating multipaths are extended to the THz frequency band. In step 11, the model described in [51] is applied to introduce sparsity into the THz channel. The detailed modifications and implementation process for THz channel simulation can be found in [137].
@@
    For RIS-assisted channel simulation, BUPTCMCCCMG-IMT2030 has modified a series of steps on the basis of the 5G standard channel model. In the RIS-assisted channel, the channel is divided into two links, Tx–RIS and RIS–Rx, and parameters are generated for them and calculated together. In step 1, the RIS is configured apart from the layout of the BS and UE. In steps 2–10, the propagation condition (LOS/NLOS), path loss, LSPs, and SSPs are generated for both Tx–RIS and RIS–Rx links. In step 11, the concatenated channel coefficients of the RIS-assisted channel are calculated by combining the channel coefficients of the Tx–RIS and RIS–Rx links and applying the radiation pattern of the RIS. For detailed modifications and implementation procedures related to RIS-assisted channel simulation, please refer to [245].
@@
    For the SAGIN channel simulation, the channel model defined in 3GPP TR 38.811 [299] is implemented and integrated into BUPTCMCCCMG-IMT2030, including modifications to the coordinate system, models, and parameters. In step 1, the ECEF coordinate system is used to locate the BS and UE. In step 2, the propagation conditions (LOS/NLOS) are assigned via a different method. In step 3, the path loss model is modified as described in VII-C. In step 4, the fast-fading parameters for the SAGIN are used, which are related to both the frequency and elevation angle.
@@
XL-MIMO
Scenario: UMi-street canyon; LOS status: LOS
Center frequency: 28 GHz; Bandwidth: 200 MHz
Position (x,y,z): BS:[0,0,3] (m); UE:[10,10,3] (m)
Antenna: BS: 256 elements ULA; UE: Single antenna
Stationary region size: 16 elements
THz
Scenario: UMi-street canyon; LOS status: LOS
Center frequency: 132 GHz; Bandwidth: 1.2 GHz
Position (x,y,z): BS:[0,0,3] (m); UE:[10,10,3] (m)
Antenna: BS: Single antenna; UE: Single antenna
Intracluster K-factor: 17.98 dB
@@
RIS size: $\times32$ elements
@@
ASA: $10^{\circ}$, $5^{\circ}$, and $1^{\circ}$]
@@
VIII-C3 Module III: Analysis
In Module III, a simple analysis of numerical and graphical results based on the generated channel coefficients is provided, including the delay spread, angular spread, reference signal receive power (RSRP), Gini index (used for sparsity analysis), network layout, and CIR of a single link.
@@ VIII-D Applications
The E-GBSM model is based on modeling using measured data, and further utilizes this model, combined with parameters obtained through fitting test data, to construct a simulator. For detailed comparative analysis, please refer to references [311, 312, 313, 137, 245]. Owing to space constraints, this paper will not elaborate further. The BUPTCMCCCMG-IMT2030 channel model has undergone rigorous calibration and validation against 3GPP TR 38.901, including key channel parameters such as ASA (azimuth spread of arrival), ASD (azimuth spread of departure), ZSA (zenith spread of arrival), ZSD (zenith spread of departure), and singular values. As demonstrated in Fig. 34, the calibration results (data 1) are in excellent agreement with the measurement data from 18 industrial partners. The 6G technology extensions are developed on the basis of the TR 38.901 framework as the baseline, ensuring fundamental correctness. Through comprehensive analysis of real-world measurement data, we have incorporated newly observed channel characteristics and enhanced modeling methodologies to extend the framework for 6G applications. This approach maintains backward compatibility with existing standards while enabling accurate characterization of emerging propagation phenomena.
Figure 34: Comparison results with the calibration of 3GPP TR 38.901.
In this section, some application examples of the typical 6G technologies (ISAC, XL-MIMO, THz, RIS, and SAGIN) are provided. The main simulation configurations for different 6G cases are listed in Table VIII.
VIII-D1 6G Feature 1–ISAC
Figure 35: Simulation results of the SD ($\text{SD}_{\text{s}}$) in the sensing channels under different numbers of shared clusters ($N_{0}$) when the total number of clusters ($N$) is 15.
Using the configurations in Table VIII, we can generate the shared and nonshared clusters of the proposed ISAC channel model in (7). The CDF of the SD metric in the sensing channels, which is the PR of the shared clusters ($N_{0}$) to the total clusters ($N$) in the ISAC channel, is shown in Fig. 35 under different numbers of shared clusters. A higher SD indicates greater potential to realize the mutual auxiliary functions of communication and sensing. As $N_{0}$ increases, $\text{SD}_{\text{s}}$ also increases. By controlling the appropriate $N_{0}$ of the proposed channel model, the desired SD value can be achieved, facilitating flexible simulation implementation of ISAC channels.
VIII-D2 6G Feature 2–XL-MIMO
On the basis of the configurations in Table VIII, the absolute cross-correlation coefficient of CIRs between the reference element and other elements in comparison with a channel generated via the standard channel model is shown in Fig. 36. With increasing spacing between two elements, the cross-correlation coefficient of the standard channel model remains stable, but that of the XL-MIMO channel continues to decrease, indicating that the correlation of the two elements becomes weaker. In addition, the appearance or disappearance of clusters and the evolution of the phase also occur. The appearance or disappearance of clusters results in a sudden change in the cross-correlation coefficients, including an increase in the appearance and a decrease in the disappearance. When no cluster appears or disappears, the evolution of the phase results in a change in the cross-correlation coefficients.
VIII-D3 6G Feature 3–THz Frequency Band
@@
Using the configurations in Table VIII, the CIR results of the generated channel at 132 GHz in the THz band were compared with those at 28 GHz in the mmWave band, as shown in Fig. 37. The THz channel has 8 clusters, and the mmWave channel has 16 clusters. Fewer multipaths are present in the THz band than in the mmWave band, which indicates that the THz channel has a sparser structure than the mmWave channel does. In addition, the delay spread in the THz channel (4.27 ns) is much lower than that in the mmWave channel (30.94 ns), according to Fig. 37. The multipaths of the THz channel are concentrated in a narrower delay range, which means that the multipaths in the THz channel are concentrated within a small time period.
@@
Using the configuration in Table VIII, the impact of channel angular spread on the performance of RIS-assisted communication is illustrated in Fig. 38. With the increase in ASA in the Tx-RIS channel, the performance of the RIS concatenated communication link will worsen. This is because the greater the angular spread, the lower the energy in the specified incident direction. Additionally, the impact of the angle-dependent characteristics of RIS reflection coefficients on RIS-assisted communication, as mentioned in Section VI-B1, has been considered in the simulation. In the legends, ”Non-ideal modulation” represents the simulation results when considering this angle-dependent characteristic. It can be observed that the non-ideal modulation results in approximately 0.5, 0.7, and 1 dB of attenuation under the three angle spreads.
@@
Using the configurations in Table VIII, the path loss of SG channels with different channel configurations is illustrated in Fig. 39. The path loss of SG channels is affected by the height of the satellites, the elevation angle of the link, and the center frequencies. An increase in satellite height results in an increase in propagation distance, which increases path loss. Furthermore, the path loss increases when a higher frequency band is used. Additionally, the elevation angle of the link can affect path loss in SG scenarios; when the elevation angle increases, path loss increases.